Find the electric potential of a point located at a distance of 0.8 m from a charge of q=8µC


1
Expert's answer
2021-10-28T17:37:53-0400

Electric potential

"V=\frac{kQ}{r}"

"V=\frac{9\times10^9\times8\times10^{-6}}{0.8}=9\times10^{4}V"